OpenStudy (anonymous):

Can someone please show me step by step how to solve this Absolute Value Equation? | 5x + 7 | = 3

OpenStudy (anonymous):

Square both sides? (5x+7)^2=3^2 25x^2+49+2(5x)(7)=9

OpenStudy (anonymous):

25x^2+70x+40=0

OpenStudy (anonymous):

5x^2+35x+8=0

OpenStudy (anonymous):

ohhh I don't think it works

OpenStudy (anonymous):

nless you try the formula

OpenStudy (anonymous):

What formula is that?

OpenStudy (anonymous):

x=-b

OpenStudy (anonymous):

sorry not thAat :D

OpenStudy (anonymous):

lol ... I'm confused but it's ok

OpenStudy (anonymous):

if you fill the quadratic equatioin into that it should work out?

OpenStudy (anonymous):

\[\frac{ -7\pm \sqrt{7^{2}- (4)(5)(3)} }{ 2a }\] like that?

OpenStudy (texaschic101):

| 5x + 7 | = 3 (5x + 7) = 3 and -(5x + 7) = 3 5x = 3 - 7 -5x - 7 = 3 5x = -4 -5x = 3 + 7 x = -4/5 -5x = 10 x = -5 x = -4/5 and x = -5

OpenStudy (texaschic101):

this should help.. http://www.purplemath.com/modules/solveabs.htm

OpenStudy (anonymous):

Thank you both... I appreciate your assistance. c:

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

Sorry I confused you!

OpenStudy (anonymous):

@yaysocks Noooo it's ok thank you for attempting. T: